OpenStudy (amonoconnor):

2(x^2 + y^2)^2 = 25(x^2 - y^2) Find points on the lemniscate (shape of graph) where the tangent is horizontal. Any and all help is greatly appreciated!

OpenStudy (freckles):

main objective is to find (x,y) when y'=0

OpenStudy (amonoconnor):

I found: \[y' = \frac{8x^3 +8xy^2-50x}{8y^3-8x^2y-50y}\]

OpenStudy (amonoconnor):

...so do I set that equal to zero and solve for either x or y?

OpenStudy (amonoconnor):

...so do I set that equal to zero and solve for either x or y?

OpenStudy (freckles):

I'm getting a little bit of sign differences in mine

OpenStudy (freckles):

\[y' = \frac{8x^3 +8xy^2-50x}{\color{red}{-}8y^3-8x^2y-50y}\] is the only difference I'm seeing in my y'

OpenStudy (amonoconnor):

Yep! Goddamn, my typos...

OpenStudy (freckles):

anyways yeah we want to find (x,y) when y'=0 and this might involved using the original equation also we can factor out 2x on top and -2y on bottom this will give us: \[y'=- \frac{x}{y} \frac{4x^2+4y^2-25}{4x^2+4y^2+25} \\ \text{ so we want } x=0 \text{ or } 4x^2+4y^2-25=0\]

OpenStudy (freckles):

and on yeah y cannot be 0 otherwise the bottom is 0

OpenStudy (freckles):

so you replace the x's in the initial equation with 0 and solve for y

OpenStudy (freckles):

we gotta also figure out how to deal with the second equation

OpenStudy (freckles):

\[4x^2+4y^2=25 \\ x^2+y^2=\frac{25}{4} \]

OpenStudy (freckles):

we can replace x^2+y^2 in the initial equation with 25/4 ...

OpenStudy (freckles):

\[2(\frac{25}{4})^2=25(x^2-y^2) \\ \frac{2}{25}(\frac{25}{4})^2=x^2-y^2 \\ \frac{2(25)}{4^2}=x^2-y^2 \\ \frac{25}{8}=x^2-y^2 \]

OpenStudy (freckles):

I guess this gives a system of equations to solve

OpenStudy (freckles):

\[\frac{25}{4}=x^2+y^2 \\ \frac{25}{8}=x^2-y^2 \]

OpenStudy (amonoconnor):

Here is the graph & source of the original equation.

OpenStudy (freckles):

you should wind up with three points

OpenStudy (freckles):

what does the (3,1) mean ?

OpenStudy (amonoconnor):

That was the point I had to find a tangent line for, in problem 31. This is #43, and asking just for horizontal tangents, but it refers to the "lemniscate" function from #31, so I thought you might want to look at it.

OpenStudy (freckles):

oh okay did you see how I got the system of equations above?

OpenStudy (amonoconnor):

Working on it:)

OpenStudy (freckles):

you might want to check my arithmetic when looking at it

OpenStudy (amonoconnor):

by "replace x's in original equation", you mean 2(x^2 + y^2)^2 = 25(x^2 - y^2)

ganeshie8 (ganeshie8):

just for practice, you could also try changing to polar coordinates. see how simple the curve looks in polar : \[2r^2 = 25\cos(2\theta)\]

OpenStudy (freckles):

yes.

OpenStudy (amonoconnor):

2/25 = y^(-2) ?

jimthompson5910 (jim_thompson5910):

@ganeshie8 r^2 = x^2+y^2, so the LHS should be 2(r^2)^2 which becomes 2r^4

jimthompson5910 (jim_thompson5910):

@freckles your arithmetic looks good. I confirmed it with geogebra

OpenStudy (freckles):

\[\frac{25}{4}=x^2+y^2 \\ \frac{25}{8}=x^2-y^2 \\ \text{ gives } 2x^2=\frac{75}{8} \text{ by adding equations }\]

OpenStudy (freckles):

oh wait are you trying to find what y is from x=0 @amonoconnor ?

ganeshie8 (ganeshie8):

Ahh good catch @jim_thompson5910

OpenStudy (freckles):

thanks @jim_thompson5910

OpenStudy (amonoconnor):

Yeah, is that wrong?

OpenStudy (freckles):

\[x=0 \\ 2(x^2+y^2)^2=25(x^2-y^2) \\ 2(0^2+y^2)^2=25(0^2-y^2) \\ 2y^4=-25y^2 \\ 2y^4+25y^2=0 \\ y^2(2y^2+25)=0 \\ y^2=0 \text{ and } 2y^2+25 \neq 0 \\ y=0 \\ (0,0) \text{ makes } y'=0\]

OpenStudy (freckles):

but

OpenStudy (freckles):

we said y cannot be 0

OpenStudy (freckles):

so (0,0) is not going to give us y'=0

OpenStudy (freckles):

so we will only have two points it looks like

OpenStudy (freckles):

or 4 lol

OpenStudy (amonoconnor):

OKay... I think I follow. I get why y cannot equal zero, but I guess I don't even know why we plugged in 0 for x... and yes, there would have to be 4 with a horizontal figure eight shape, right?

OpenStudy (freckles):

because y'=0 when the numerator equals 0 as long as it doesn't also make the bottom 0

OpenStudy (amonoconnor):

Ahh... no rise, over some run. Slope of 0. Gotcha:)

OpenStudy (freckles):

so you have to consider the other factor on top=0

OpenStudy (freckles):

\[y'=- \frac{x}{y} \frac{4x^2+4y^2-25}{4x^2+4y^2+25} \\ \text{ so we want } x=0 \text{ or } 4x^2+4y^2-25=0 \] x=0 gives y=0 so (0,0) won't work we look at the other factor on top \[4x^2+4y^2-25=0 \\ 4x^2+4y^2=25 \\ x^2+y^2=\frac{25}{4}\] plug this into the original equation if you haven't already and you should get another equation in terms of x and y giving you a system to solve

OpenStudy (amonoconnor):

How does the far left chunk equal the middle?

ganeshie8 (ganeshie8):



OpenStudy (freckles):

far left chunk?

OpenStudy (freckles):

and what middle

OpenStudy (amonoconnor):

Sorry, of the 3 expressions that are separated by equals signs. 4x^2 + 4y^2 − 25 = 04x^2 + 4y^2

OpenStudy (amonoconnor):

These 2.

OpenStudy (freckles):

those aren't equal

OpenStudy (freckles):

are you talking these 3 equations? \[4x^2+4y^2-25=0 \\ 4x^2+4y^2=25 \\ x^2+y^2=\frac{25}{4} \]

OpenStudy (freckles):

the second equation I got by adding 25 on both sides

OpenStudy (amonoconnor):

Yes:) Those!

OpenStudy (freckles):

the third equation I was just showing I divided by 4 on both sides

OpenStudy (freckles):

those are three the same equation

OpenStudy (freckles):

I was just solving for x^2+y^2 because I see that I can easily replace x^2+y^2 in the original equation if I do so

OpenStudy (freckles):

do you know how got the first equation "?

OpenStudy (freckles):

that is the 4x^2+4y^2-25=0

OpenStudy (amonoconnor):

Yes, I get that... I'm good there:) And now I can plug in 25/4 for (x^2 + y^2) ?

OpenStudy (freckles):

right on

OpenStudy (freckles):

I actually did this above I will just copy and paste it here so you can check your work: \[2(\frac{25}{4})^2=25(x^2-y^2) \\ \frac{2}{25}(\frac{25}{4})^2=x^2-y^2 \\ \frac{2(25)}{4^2}=x^2-y^2 \\ \frac{25}{8}=x^2-y^2\]

OpenStudy (freckles):

You are almost there by the way... Just a little further.

OpenStudy (amonoconnor):

1/4 = (x^2 - y^2) ?

OpenStudy (amonoconnor):

Then do I solve for one variable, and plug it back in?

OpenStudy (freckles):

\[2(x^2+y^2)^2=25(x^2-y^2) \\ \text{ we have } x^2+y^2=\frac{25}{4} \\ 2(\frac{25}{4})^2=25(x^2-y^2) \\ \text{ solving for } (x^2-y^2) \text{ by dividing both sides by 25 } \\ \frac{2}{25}(\frac{25}{4})^2=x^2-y^2\] we try simplifying this left hand side one more time

OpenStudy (freckles):

try simplifying this left hand side one more time *

OpenStudy (freckles):

\[\frac{2}{25}(\frac{25}{4})^2 =\frac{2}{25}\cdot \frac{25^2}{4^2} =\frac{25^2}{25} \cdot \frac{2}{4^2}=25^{2-1} \cdot \frac{2}{16}=25^{1}\cdot \frac{1}{8}=25 \cdot \frac{1}{8}=\frac{25}{8}\]

OpenStudy (freckles):

so we have the following two equations: \[x^2+y^2=\frac{25}{4} \\ x^2-y^2=\frac{25}{8}\] I would solve the system by elimination it is already setup for eliminiantion just add equations (eliminating y allowing you to solve for x)

OpenStudy (amonoconnor):

\[2(x^2+y^2)^2 = 25(x^2-y^2)\] \[2(\frac{25}{4})^2 = 25(x^2-y^2)\] \[(\frac{50}{4})^2 = 25(x^2-y^2)\] \[(\frac{1}{25})(\frac{50}{4})^2 = (x^2-y^2)\] \[(\frac{50}{100})^2 = (x^2-y^2)\] (1/2)^2 = (x^2 - y^2) 1/4 = (x^2 - y^2)

OpenStudy (freckles):

there is a square on 25 you ignored it in your third line by choosing to multiply 25 by 2

OpenStudy (amonoconnor):

It wouldn't just translate in?

OpenStudy (freckles):

if you had 2^2 ... you could use law of exponents and invite 2 into the square club that 25 is already in

OpenStudy (freckles):

that is if you had: \[2^2(\frac{25}{4})^2=( \frac{2 \cdot 25}{4})^2 \]

OpenStudy (freckles):

but 2 cannot be in the square club because 2 does not have that 2 exponent thingy

OpenStudy (amonoconnor):

I'm just tired, you're right lols!

OpenStudy (freckles):

you did the same thing with 1/25